OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (3, 10) and (7, 8). (5, 9)?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

\[\huge \left(\frac{3 + 7}{2} , \; \; \frac{10 + 8}{2} \right) \implies (5,9)\]